OpenStudy (anonymous):

A student takes a 20 question true/false exam and guesses on each question. Find the probability of passing if the lowest passing grade is 15 correct out of 20. Would you consider this event likely to occur?

OpenStudy (anonymous):

\[\dbinom{20}{15}\times\frac{1}{2^{20}}\] not so likely

OpenStudy (anonymous):

\[\frac{15504}{1048576}\]

OpenStudy (anonymous):

.0148 rounded